Ticket Number: HEA-NOC/20060811-5 Ticket Status: UPDATE
Ticket Type: scheduled Resolver: Eircom
Ticket Opened: 20060811 16:15 UTC+1 Problem Start: 20060815 15:00 UTC+1
Ticket Update: 20060816 15:10 UTC+1 Problem End:
Site/line: CIT, DKIT, GMIT, IADT, IT Athlone, IT Blanchardstown, IT
Carlow, IT Dundalk, IT Sligo, IT Tralee, ITTallaght, ITnet,
The two STM-1s connecting ITnet in Brooklawn House are to be rehomed. The
first of these is expected to complete testing by Eircom on Monday, and
the move will take place once a successful test is complete.
This ticket will be updated with an exact time as soon as one is available
from the telco.
During the move, an outage to the sites listed above may be observed while
failover is configured and tested. Every effort will be made to avoid or
minimise the duration of this outage.
20060811 16:07 DW Inform Martin McCarrick and open ticket.
20060814 14:18 DW Eircom will perform final tests on the Citywest STM-1
tomorrow and expect handover before lunchtime. Cutover is expected
tomorrow afternoon, pending confirmation from Eircom in the morning.
20060815 12:12 DW Eircom report errors detected in the cross connect which
are now being investigated. Cutover will not take place until these have
been dealt with.
20060815 13:57 DW Eircom are retesting and currently expect to perform the
cutover at 15:00.
20060815 14:34 DW Eircom have requested a further postponement to later
today or tomorrow. Will update with more info as soon as it's available.
20060815 16:44 DW Eircom will attempt to deliver the circuit shortly - if
successful the link cutover will take place after 17:00 (so affected sites
should back up on to the alternate path) and the routing change should take
place out of hours.
20060815 17:16 DW Link cutover about to begin, with sites backing up onto
resilient path; routing change to take place after 17:30.
20060815 18:22 DW Cutover was unsuccessful. All sites backed up onto
resilient path with a few seconds convergence time. Have now scaled back
and will make another attempt tomorrow after troubleshooting.
20060816 11:57 DW Next attempt to be made at 13:30 today. As was the case
yesterday, sites will back up onto the resilient path while the switch
20060816 13:22 DW A problem with the fibre run in Esat-X prevents us from
proceeding. Awaiting a rescheduled time from Eircom.
20060816 13:58 DW Link now running clear all the way to our cage,
changeover scheduled for 14:30.
20060816 14:45 DW Cutover to new SDH link has started. All affected sites
have backed up onto the resilient path, and the new link is testing
successfully. Routing change to divert traffic onto this new link will
take place shortly.
20060816 15:08 DW Routing change has taken place. Traffic is now routing
via the new circuit. There was a 7 second outage while routing converged.
Please notify the NOC if any further problems are experienced. Ticket
remains open for monitoring.
Time to Fix:
Not yet known.
This ticket can be monitored at http://www.hea.net/tickets/20060811-5